OpenStudy (jjuden):

The equation of a circle is (x+4)^2+(y−6)^2=16. The center of the circle with the given equation is?

OpenStudy (misty1212):

HI!!

OpenStudy (jjuden):

hi

OpenStudy (misty1212):

\[9(x-h)^2+(y-k)^2=r^2\\ (x-(-4))^2+(y-6)^2=4^2\]

OpenStudy (misty1212):

ignore the 9, it is a typo!

OpenStudy (jjuden):

lol ok

OpenStudy (misty1212):

you can read off the center from the formula

OpenStudy (misty1212):

you got it?

OpenStudy (misty1212):

center is \((h,k)\) what is it in your example?

OpenStudy (jjuden):

so (-4,6)

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (jjuden):

ok thank you :)

OpenStudy (misty1212):

btw the radius is \(4\) not that you were asked

OpenStudy (jjuden):

ty :*